OpenStudy (anonymous):

use the epsilon - little delta definition to prove that the lim as x approaches 2 (6x^2 - 3) =21

OpenStudy (helder_edwin):

Prove that \[ \large \lim_{x\to2}(6x^2-3)=21. \] Let \(\varepsilon>0\) then \[ \large |6x^2-3-21|=|6x^2-24|=6|x^2-4|=6|x-2|\cdot|x+2|. \] Let \(\delta=1\) so \[ \large |x-2|<\delta=1 \] \[ \large -1<x-2<1 \] \[ \large 1<x<3 \] \[ \large 3<x+2<5\qquad\text{so}\qquad |x+2|<5 \] Then \[ \large |6x^2-24|=6|x-2|\cdot|x+2|<6\cdot5|x-2|<\varepsilon \] \[ \large \Rightarrow |x-2|<\varepsilon/30. \] Therefore \(\delta=\min\{1,\varepsilon/30\}\).

OpenStudy (anonymous):

thank you so much:3 life saver!

OpenStudy (helder_edwin):

u r welcome